The capacitors have values C1 = 2.5 μF and C2 = 4.0 μF, C3 =

6.0 μF C4 = 8.0 μF and the potential difference across the battery is

12.0 V. Assume that the capacitors are connected in series. Find the

equivalent capacitance of the circuit and solve for the potential

difference across each capacitors.


1
Expert's answer
2022-03-16T09:41:54-0400
"\frac{1}{C}=\frac{1}{C_1}+\frac{1}{C_2}+\frac{1}{C_3}+\frac{1}{C_4}""\frac{1}{C}=\frac{1}{2.5\:\rm \mu F}+\frac{1}{4.0\:\rm \mu F}+\frac{1}{6.0\:\rm \mu F}+\frac{1}{8.0\:\rm \mu F}""C=1.1\:\rm \mu F"


The charge on the plates of each capacitor

"q=CV\\
=1.1*10^{-6}*12.0=1.3*10^{-5}\:\rm C=13\:\mu C"

The potential difference between plates of capacitors

"V_1=\frac{q}{C_1}=\frac{13}{2.5}=5.2\:\rm V"

"V_2=\frac{q}{C_2}=\frac{13}{4.0}=3.2\:\rm V"

"V_3=\frac{q}{C_3}=\frac{13}{6.0}=2.1\:\rm V"

"V_4=\frac{q}{C_4}=\frac{13}{8.0}=1.6\:\rm V"